During the year, the XYZ company paid $180 in interest. The company is in the 25% tax bracket. It issued an additional $50 in notes payable and an additional $50 in long term debt. The firm paid $25 in dividends and brought back 30$ in common stock. It also increased its short term investments by $15. Based on the information above, calculate the Free Cash Flow for XYZ company.
Answer #1

Increase in working capital = increase in short term investments - Increase in notes payable  = 50- 15 = -35

Interest paid out of income net of taxes = 180 * ( 1-25%) = 135

FCF = Net income + non cash expenses - increase in working capital - capital expenditures

= 135 +0 - (-35) -0

= 170